# Self-Dual Codes and LCD Codes in Sum-Rank Metric

### v1

@@
Abstract: Sum-rank codes are an important class of codes which can be utilized for linear network coding, space-time coding and distributed storage. Based on the duality theory of sum-rank codes [Byrne, Gluesing-Luerssen, Ravagnani, IEEE TIT, 2021], it is interesting to study self-dual sum-rank codes and linear complementary dual (LCD) sum-rank codes. Firstly, we characterize the dual codes of some sum-rank codes. Then we define self-dual sum-rank codes and LCD sum-rank codes, provide some basic properties of such codes and then obtain two methods of constructing self-dual sum-rank codes and LCD sum-rank codes from Euclidean self-dual codes and Euclidean LCD codes. Some particular examples especially some cyclic self-dual sum-rank codes and cyclic LCD sum-rank codes with good parameters are also provided. At last, we prove that there exist asymptotically good self-dual sum-rank codes.
Index Terms: sum-rank code; Euclidean self-dual code; Euclidean LCD code; cyclic code.
@@ 1 Introduction
Self-dual codes are a fascinating family of codes which has inspired many researchers and has seen a tremendous growth in the past 50 years. They have many applications in coding theory, cryptography and combinatorics, such as [17, 56, 61]. The earlier results about self-dual codes over small fields can be seen in [18, 19, 20, 21] and [32, Chapter 9]. The theory for the classification of self-dual codes was provided in [62]. There are many studies on constructing cyclic self-dual codes in [14, 68, 16]. Some researchers used algebraic technique to find good self-dual codes including [43, 46, 53]. Also, there have been many papers, which constructed self-dual codes by using GRS codes, see [22, 29, 31, 34, 38, 69]. The minimum distance of a binary self-dual code $C$ of length $n$ satisfies
$d_{H}(C)\leq 4\lfloor\frac{n}{24}\rfloor+6$
if $n\equiv 22\,(\textrm{mod}\,24),$ and
$d_{H}(C)\leq 4\lfloor\frac{n}{24}\rfloor+4$
otherwise. A binary self-dual code attaining these upper bound is called an extremal self-dual code [47, Chapter 19]. We refer the reader to [4, 21, 25] for the construction of extremal or optimal binary self-dual codes. The authors of [24] got the result that the minimum distance of a self-dual code $C$ over $\mathbb{F}_{4}$ of length $n$ satisfies
@@
Apart from Hamming metric, some researchers turned to self-dual codes in other metrics, because in some cases, different metrics lead to different alphabets, different inner products as well as different applications. In [27, 64], the MacWilliams identities for codes in rank metric was researched. According to the MacWilliams identities, the inherent trade-off between dimension and minimum distance for a code is reversed for its dual code: if a code has high dimension and low minimum distance, then its dual code will have low dimension and high minimum distance, so self-dual rank-metric codes raised some researchers’ interest. In [26], Galvez and Kim obtained self-dual rank-metric codes from a self-dual rank-metric code of smaller size using a method called the building-up construction. In [26, 57], the self-dual rank-metric codes of small sizes over small finite fields were classified. In [58], Nebe and Willems focused on the self-dual MRD codes (MRD codes are those rank-metric codes attaining Singleton-type bound).
@@
LCD codes are another interesting family of codes related to the duality theory which was introduced by Massey [55] and have been investigated deeply. They have applications in symmetric cryptosystems [6], prevention from hardware Trojan horse insertions [10] and construction of entanglement-assisted quantum error-correcting codes [30]. There are many references for the construction of LCD codes, such as [11, 54, 55]. Many LCD BCH codes with very good parameters were constructed in [40, 41]. The hull variation problem was proposed and studied in our previous paper [13]. The monomial equivalence of Galois LCD codes was studied in [46].
@@
Apart from the characterization and construction of LCD codes in Hamming metric, LCD codes in various metrics were considered by some researchers. The authors of [37] found that a special class of rank-metric codes named Gabidulin codes over $\mathbb{F}_{2^{m}}$ of full length $m$ generated by the trace-orthogonal-generator matrices were shown to be both LCD codes and MRD codes. Motivated by this, some other researchers became interested in LCD codes in rank-metric. In [39], more LCD-MRD codes were constructed. Liu and Liu [45] pointed out the relationship between LCD rank-metric codes and Euclidean LCD codes. Based on the relationship, they also constructed some LCD-MRD codes and gave an application of rank-metric LCD codes in decoding algorithm. In [63], the authors applied a class of LCD rank-metric codes in construction of $\mathcal{T}$-direct codes which were employed for the $\mathcal{T}$-user F-Adder channel. The authors of [70] found that one can always construct LCD generalized Gabidulin codes under some conditions.
@@
Rank-metric codes are used to achieve a reliable communication over multiplicative-additive finite field matrix channels. The authors of [59] considered blocks of $n$ consecutive uses of the channel, so they obtained sum-rank codes which can not only be used to reduce the size of the network alphabet but also correct more errors. Thus, to some extent, the sum-rank codes are a generalization or improved version of rank-metric codes and have similar applications in theory and practice. Therefore, similar to the cases in rank metric, we consider self-dual codes and LCD codes in sum-rank metric based on the duality theory of sum-rank codes. In this paper, our main contributions are as follows:
  • In Lemma 3.1, Lemma 3.6 and Corollary 3.7, we characterize the dual codes of two classes of sum-rank codes.
@@
  • In Theorem 4.4 and Theorem 4.16, we provide two methods of constructing self-dual codes and LCD codes in sum-rank metric from Euclidean self-dual codes and Euclidean LCD codes.
  • Some cyclic self-dual codes in sum-rank metric are constructed in Theorem 4.9, Theorem 4.22 and so on. Some cyclic LCD codes in sum-rank metric are constructed in Proposition 4.15, Proposition 4.25 and so on.
@@
This paper is organized as follows. In Section 2, some necessary knowledge about Euclidean self-dual codes, Euclidean LCD codes, cyclic codes and sum-rank codes are presented. In Section 3, we give the characterization of the dual codes of some sum-rank codes. In Section 4, we define self-dual sum-rank codes and LCD sum-rank codes and give some properties of self-dual sum-rank codes. Then we introduce two methods of constructing self-dual sum-rank codes and LCD sum-rank codes from Euclidean self-dual codes and Euclidean LCD codes and provide some particular examples with good parameters and algebraic structures. In Section 5, we give an asymptotically good sequence of self-dual sum-rank codes. In Section 6, we conclude this paper.
@@ 2 Preliminaries
In this section, we provide some basic knowledge about Euclidean self-dual codes and Euclidean LCD codes, cyclic codes as well as sum-rank codes which will be used later. Throughout the paper, $\mathbb{F}_{q}$ is the finite field with $q$ elements where $q$ is a power of a prime. Let $\bm{x}=(x_{1},\ldots,x_{n})$, $\bm{y}=(y_{1},\ldots,y_{n})\in\mathbb{F}_{q}^{n}$, the Hamming distance between $\bm{x}$ and $\bm{y}$ is
@@
Let $C\subseteq\mathbb{F}_{q}^{n}$ be a code, its minimum Hamming distance is as follows:
@@
If $C$ is a linear subspace of $\mathbb{F}_{q}^{n}$, $C$ is called a linear code. A linear code $C$ in $\mathbb{F}_{q}^{n}$ with dimension $k$ and minimum Hamming distance $d$ is called an $[n,k,d]_{q}$-linear code. A code sequence $C_{1}$, $C_{2}$, $\cdots$ is said to be asymptotically good if the length $n_{i}$ of $C_{i}$ goes to infinity and for $i=1,2,\ldots$, both the rate $R_{H}(C_{i})$ and the relative minimum Hamming distance $\Delta_{H}(C_{i})$ are positively bounded from below.
@@ 2.1 Euclidean Self-Dual Codes and Euclidean LCD Codes
We firstly review some basic notions and results about Euclidean self-dual codes and Euclidean LCD codes, although most of them are well-known. Let $\bm{x}=(x_{1},\ldots,x_{n})$, $\bm{y}=(y_{1},\ldots,y_{n})\in\mathbb{F}_{q}^{n}$. Then the Euclidean inner product of $\bm{x}$ and $\bm{y}$ is defined as
@@
Let $C$ be a linear code in in $\mathbb{F}_{q}^{n}$. Then the dual code of $C$ is
@@
The lemma below is an upper bound on the minimum Hamming distance of a Euclidean self-dual code over $\mathbb{F}_{4}$ from [24]. In our work, we will use the upper bound to obtain an upper bound on the minimum sum-rank distance of self-dual sum-rank codes.
@@
([24, Corollary 3.4]) Let $C$ be a Euclidean self-dual code over $\mathbb{F}_{4}$ of length $n$. Then $d_{H}(C)\leq 4\lfloor\frac{n}{12}\rfloor+4$.
Next, we review a sequence of asymptotically good self-dual codes in [43]. Let $G$ be an abelian group of order $n$. By $\mathbb{F}_{q}G$ we denote the group algebra, i.e., the $\mathbb{F}_{q}$-vector space with basis $G=\{x_{1},\ldots,x_{n}\}$ and with the multiplication induced by the group multiplication of $G$. Any vector $\bm{a}=(a_{1},\ldots,a_{n})\in\mathbb{F}_{q}^{n}$ can be identified with an element $\sum\limits_{i=1}^{n}a_{i}x_{i}\in\mathbb{F}_{q}G$. An ideal $C$ of $\mathbb{F}_{q}G$ is called an $\mathbb{F}_{q}G$-code. Let $(\mathbb{F}_{q}G)^{2}:=\mathbb{F}_{q}G\times\mathbb{F}_{q}G=\{(a,b)\,|\,a,b\in\mathbb{F}_{q}G\}$ which is an $\mathbb{F}_{q}G$-module. Any $\mathbb{F}_{q}G$-submodule of $(\mathbb{F}_{q}G)^{2}$ is called a 2-quasi-$\mathbb{F}_{q}G$ code. In addition, we define the $q$-cyclotomic cosets $C_{i}^{(q,n)}=\{i,iq,\ldots,iq^{l-1}\}{\pmod{n}}$, $l$ is the smallest positive integer such that $iq^{l}\equiv i{\pmod{n}}$. $\{0,1,\ldots,n-1\}$ is partitioned into several $q$-cyclotomic cosets. Obviously, $\{0\}$ is a $q$-cyclotomic coset which we call the trivial $q$-cyclotomic coset. By $\mu_{q}(n)$ we denote the minimal size of the non-trivial $q$-cyclotomic cosets. Throughout the paper, we keep these notations.
Lemma 2.3.
([43, Theorem 1.1]) Let $n_{1}$, $n_{2}$, $\ldots$ be odd positive integers coprime to $q$ satisfying that $n_{i}$ goes to infinity with $i$ and $\lim\limits_{i\rightarrow\infty}\frac{\log_{q}n_{i}}{\mu_{q}(n_{i})}=0$. For $i=1,2,\ldots$, let $G_{i}$ be any abelian group of order $n_{i}$. If $-1$ is a square in $\mathbb{F}_{q}$, then there exist self-dual $2$-quasi-$\mathbb{F}_{q}G_{i}$ codes $C_{i}$, $i=1,2,\ldots$, such that the code sequence $C_{1}$, $C_{2}$, $\ldots$ is asymptotically good.
By [43], there exist odd integers $n_{1}$, $n_{2}$, $\ldots$ coprime to $q$ satisfying that $\lim\limits_{i\rightarrow\infty}\frac{\log_{q}n_{i}}{\mu_{q}(n_{i})}=0$. Hence, if $-1$ is a square in $\mathbb{F}_{q}$, there exists a sequence of asymptotically good self-dual codes over $\mathbb{F}_{q}$.
We end up this subsection with a famous and important theorem about Euclidean LCD codes from [11].
@@
([11, Corollary 5.4]) Let $q$ be a power of a prime with $q>3$. Then, an $[n,k,d]_{q}$-linear Euclidean LCD code exists if there is an $[n,k,d]_{q}$-linear code.
However, when the length $n$ and dimension $k$ are certain, in some cases, we cannot determine the exact value of $d_{LCD}(n,k)_{2}=\max\{d_{H}(C)\,:\,\text{$C$ is an $[n,k,d_{H}(C)]_{2}$ LCD code}\}$ and $d_{LCD}(n,k)_{3}=\max\{d_{H}(C)\,:\,\text{$C$ is an $[n,k,d_{H}(C)]_{3}$ LCD code}\}$.
@@ 2.2 Cyclic Codes
A linear code $C\subseteq\mathbb{F}_{q}^{n}$ is said to be cyclic if $(c_{0},c_{1},\ldots,c_{n-1})\in C$ implies $(c_{n-1},c_{0},\ldots,c_{n-2})\\ \in C$. In the subsection, we recall some facts about cyclic codes especially those results about cyclic self-dual codes and cyclic LCD codes.
By identifying $(c_{0},c_{1},\ldots,c_{n-1})\in\mathbb{F}_{q}^{n}$ with $c_{0}+c_{1}x+c_{2}x^{2}+\cdots+c_{n-1}x^{n-1}\in\mathbb{F}_{q}[x]/\langle x^{n}-1\rangle$, the corresponding cyclic code $C\subseteq\mathbb{F}_{q}^{n}$ is an ideal of $\mathbb{F}_{q}[x]/\langle x^{n}-1\rangle$. Since every ideal of $\mathbb{F}_{q}[x]/\langle x^{n}-1\rangle$ is principal, there exists a monic polynomial $g(x)$ of the smallest degree such that $C=\langle g(x)\rangle$ where $g(x)\,|\,x^{n}-1$. In addition, $g(x)$ is unique and called the generator polynomial.
Let $n$ be a positive integer with ${\rm gcd}(n,q)=1$ and $m$ be the smallest positive integer such that $q^{m}\equiv 1{\pmod{n}}$. Let $\alpha$ be a generator of $\mathbb{F}_{q^{m}}^{*}$ and put $\beta=\alpha^{\frac{q^{m}-1}{n}}$. Then $\beta$ is a primitive $n$-th root of unity. For $10\leq i\leq n-$ 1, let $m_{i}(x)$ denote the minimal polynomial of $\beta^{i}$ over $\mathbb{F}_{q}$. Then
$x^{n}-1={\rm lcm}(m_{0}(x),m_{1}(x),\ldots,m_{n-1}(x))\text{ and }m_{i}(x)=\prod\limits_{j\in C_{i}^{(q,n)}}(x-\beta^{j})$
where the $q$-cyclotomic-coset $C_{i}^{(q,n)}$ is defined as previously mentioned. Note that we use $i{\pmod{n}}$ to denote the unique integer in the set $\{0,1,\ldots,n-1\}$, which is congruent to $i$ modulo $n$. Thus, we have $m_{i}(x):=m_{i{\pmod{n}}}(x)$.
@@
Let $C_{(q,n,\delta,b)}$ be the cyclic code of length $n$ with generator polynomial $g_{(q,n,\delta,b)}(x)$. Then $C_{(q,n,\delta,b)}$ is called a BCH code with designed distance $\delta$.
BCH codes are an important class of cyclic codes. It is well-known that the minimum distance of $C_{(q,n,\delta,b)}$ is greater or equal to $\delta$, see [32].
There are many studies on constructing cyclic self-dual codes and cyclic LCD codes, such as [14], [23], [33], [36], [40] and [41]. To meet our needs, we focus on those cases where cyclic self-dual codes and cyclic LCD codes over $\mathbb{F}_{4}$ are considered. In the following, we present some results about them which will be used later. We begin with a family of cyclic self-dual codes.
Lemma 2.6.
([14, Theorem 4.1]) Let $n=\frac{2(2^{sm}-1)}{u}$, where $m=3,5,7,\ldots$ and $u$ is a divisor of $12^{sm}-$ 1. Then there exists a family of Euclidean self-dual repeated-root cyclic codes over $\mathbb{F}_{2^{s}}$ with the length $n$ and the minimum distance at least $\sqrt{\frac{2^{s-1}n}{u}}-\frac{2^{s}}{u}$.
We know that a cyclic code of length $n$ over $\mathbb{F}_{q}$ is an LCD code if $-1$ is a power of $q$ modulo $n$ (see [40]), so cyclic codes over $\mathbb{F}_{q}$ with length $n=\frac{q^{m}+1}{N}$ are always LCD codes, where $11\leq N\leq q^{m}+$ 1 is a divisor of $q^{m}+1$. There are many cyclic LCD codes over $\mathbb{F}_{q}$ with length $n=\frac{q^{m}+1}{N}$. The following is an example from [23].
Lemma 2.7.
([23, Theorem 3.4]) Suppose $n=\frac{q^{m}+1}{q+1}$, where $m>5$ is odd. For $\delta=lq^{\frac{m-1}{2}}+1$ with $12\leq l\leq q-$ 1, the code $C_{(q,n,\delta,1)}$ and $C_{(q,n,\delta+1,0)}$ have parameters $[n,k,\geq\delta]$ and $[n,k-1,\geq 2\delta]$, respectively, where

### v2

@@
Abstract: Sum-rank codes are an important class of codes which can be utilized for linear network coding, space-time coding and distributed storage. They can not only reduce the size of network alphabet but also detect and correct more errors. Based on the duality theory of sum-rank codes [Byrne, Gluesing-Luerssen, Ravagnani, IEEE TIT, 2021] and those related theory of rank-metric codes, it is significant to study self-dual codes and linear complementary dual (LCD) codes in sum-rank metric. In this paper, we introduce the notion of self-dual codes and LCD codes in sum-rank metric, and obtain two methods of constructing self-dual sum-rank codes and LCD sum-rank codes from Euclidean self-dual codes and Euclidean LCD codes. Some examples of cyclic self-dual sum-rank codes and cyclic LCD sum-rank codes with good parameters are provided. In addition, we prove that there exist asymptotically good self-dual sum-rank codes.
Keywords: sum-rank code; Euclidean self-dual code; Euclidean LCD code; cyclic code.
@@ 1 Introduction
In classical coding theory, self-dual codes are a fascinating family of codes which has inspired many researchers and has seen a tremendous growth in the past 50 years. They have many applications in coding theory, cryptography and combinatorics. The earlier results about self-dual codes over small fields can be seen in [20, Chapter 9] and [47]. Many researchers tried to construct self-dual codes by using GRS codes and algebraic techniques, see [14, 22, 26, 31, 32, 40, 43, 53]. Also, there are some studies on constructing cyclic self-dual codes, such as [11, 52, 13]. The authors of [16] got the result that the minimum distance of a self-dual code $C$ over $\mathbb{F}_{4}$ of length $n$ satisfies
@@
LCD codes are another interesting family of codes related to the duality theory which was introduced by Massey [42] and have been investigated deeply. They have applications in symmetric cryptosystems and constructions of entanglement-assisted quantum error-correcting codes. There are many references for the constructions of LCD codes, such as [9, 41, 42]. Many LCD BCH codes with very good parameters were constructed in [28, 29].
@@
Apart from codes with Hamming metric, some researchers turned to self-dual codes and LCD codes in other metrics. Rank-metric codes are used to achieve a reliable communication over multiplicative-additive finite field matrix channels. In [18, 48], the MacWilliams identities for codes in rank metric were provided which motivated researchers to consider self-dual codes and LCD codes in rank metric. In [17], Galvez and Kim obtained self-dual rank-metric codes from a self-dual rank-metric code of smaller size using a method called the building-up construction. In [17, 44], the self-dual rank-metric codes of small sizes over small finite fields were classified. In [45], Nebe and Willems focused on the self-dual MRD codes (MRD codes are those rank-metric codes attaining Singleton-type bound). As for LCD rank-metric codes, the authors of [25] found that a special class of rank-metric codes named Gabidulin codes over $\mathbb{F}_{2^{m}}$ of full length $m$ generated by the trace-orthogonal-generator matrices were shown to be both LCD codes and MRD codes. In [27], more LCD-MRD codes were constructed. Liu and Liu pointed out the relationship between LCD rank-metric codes and Euclidean LCD codes in [34]. Based on the relationship, they also constructed some LCD-MRD codes and gave an application of rank-metric LCD codes in decoding algorithm. The authors of [54] found that one can always construct LCD generalized Gabidulin codes under some conditions.
@@
Sum-rank codes are a generalization or improved version of rank-metric codes and have important applications in theory and practice. In [46], the authors considered blocks of $n$ consecutive uses of the matrix channel and then introduced the model of sum-rank codes. Compared with rank-metric codes, sum-rank codes can not only reduce the size of network alphabet but also detect and correct more errors. Therefore, it is significant to research self-dual codes and LCD codes in sum-rank metric based on the duality theory of sum-rank codes and those related theory of rank-metric codes. In this paper, our main contributions are as follows:
  • In Theorems 3.4, 3.8 and Corollary 3.9, we characterize the dual codes of two classes of sum-rank codes.
@@
  • In Theorems 4.4 and 4.19, we provide two methods of constructing self-dual codes and LCD codes in sum-rank metric from Euclidean self-dual codes and Euclidean LCD codes. Using these methods, cyclic self-dual codes and cyclic LCD codes in sum-rank metric are constructed in Theorems 4.10, 4.25, Propositions 4.18, 4.29 and so on. Some of them have good parameters which achieve the upper bound in Theorem 3.3.
@@
This paper is organized as follows. In Section 2, some necessary knowledge about Euclidean self-dual codes, Euclidean LCD codes, cyclic codes and sum-rank codes are presented. In Section 3, we characterize the dual codes of some sum-rank codes. In Section 4, we define self-dual sum-rank codes and LCD sum-rank codes and give some properties of self-dual sum-rank codes. Then we introduce two methods of constructing self-dual codes and LCD codes in sum-rank metric from Euclidean self-dual codes and Euclidean LCD codes and also provide several examples with good parameters and algebraic structures. In Section 5, we give an asymptotically good sequence of self-dual sum-rank codes. In Section 6, we conclude this paper.
@@ 2 Preliminaries
In this section, we provide some basic knowledge about Euclidean self-dual codes, Euclidean LCD codes, cyclic codes as well as sum-rank codes which will be used later. Throughout this paper, $\mathbb{F}_{q}$ is the finite field with $q$ elements where $q$ is a power of a prime. Let $\bm{x}=(x_{1},\ldots,x_{n})$, $\bm{y}=(y_{1},\ldots,y_{n})\in\mathbb{F}_{q}^{n}$, the Hamming distance between $\bm{x}$ and $\bm{y}$ is
@@
Let $C\subseteq\mathbb{F}_{q}^{n}$ be a code, its minimum Hamming distance is as follow:
@@
If $C$ is a linear subspace of $\mathbb{F}_{q}^{n}$, $C$ is called a linear code. A linear code $C$ in $\mathbb{F}_{q}^{n}$ with dimension $k$ and minimum Hamming distance $d$ is called an $n,k,d]_{q}-$linear code. A code sequence $C_{1}$, $C_{2}$, $\cdots$ is said to be asymptotically good if the length $n_{i}$ of $C_{i}$ goes to infinity and for $i=1,2,\ldots$, both the rate $R_{H}(C_{i})$ and the relative minimum Hamming distance $\Delta_{H}(C_{i})$ are positively bounded from below.
@@ 2.1 Euclidean Self-Dual Codes and Euclidean LCD Codes
We firstly review the definitions and some important results about Euclidean self-dual codes and Euclidean LCD codes. Let $\bm{x}=(x_{1},\ldots,x_{n})$, $\bm{y}=(y_{1},\ldots,y_{n})\in\mathbb{F}_{q}^{n}$. Then the Euclidean inner product of $\bm{x}$ and $\bm{y}$ is defined as
@@
Let $C$ be a linear code in $\mathbb{F}_{q}^{n}$. Then the dual code of $C$ is
@@
The lemma below is an upper bound on the minimum Hamming distance of a Euclidean self-dual code over $\mathbb{F}_{4}$ from [16]. In our work, we will derive an upper bound on the minimum sum-rank distance of self-dual sum-rank codes from it.
@@
([16, Corollary 3.4]) Let $C$ be a Euclidean self-dual code over $\mathbb{F}_{4}$ of length $n$. Then $d_{H}(C)\leq 4\lfloor\frac{n}{12}\rfloor+4$.
We end up this subsection with a famous theorem about Euclidean LCD codes from [9].
@@
([9, Corollary 5.4]) Let $q$ be a power of a prime with $q>3$. Then, an $n,k,d]_{q}-$linear Euclidean LCD code exists if there is an $n,k,d]_{q}-$linear code.
@@ 2.2 Cyclic Codes
A linear code $C\subseteq\mathbb{F}_{q}^{n}$ is said to be cyclic if $(c_{0},c_{1},\ldots,c_{n-1})\in C$ implies $(c_{n-1},c_{0},\ldots,c_{n-2})\\ \in C$. In the subsection, we recall some facts about cyclic codes.
By identifying $(c_{0},c_{1},\ldots,c_{n-1})\in\mathbb{F}_{q}^{n}$ with $c_{0}+c_{1}x+c_{2}x^{2}+\cdots+c_{n-1}x^{n-1}\in\mathbb{F}_{q}[x]/\langle x^{n}-1\rangle$, the corresponding cyclic code $C\subseteq\mathbb{F}_{q}^{n}$ is an ideal of $\mathbb{F}_{q}[x]/\langle x^{n}-1\rangle$. Since every ideal of $\mathbb{F}_{q}[x]/\langle x^{n}-1\rangle$ is principal, there exists a monic polynomial $g(x)$ of the smallest degree such that $C=\langle g(x)\rangle$ where $|\,x^{n}-1. g(x)g(x)$ is unique and called the generator polynomial of $C$.
Let $n$ be a positive integer with ${\rm gcd}(n,q)=1$ and $m$ be the smallest positive integer such that $q^{m}\equiv 1{\pmod{n}}$. Let $\alpha$ denote a generator of $\mathbb{F}_{q^{m}}^{*}$ and put $\beta=\alpha^{\frac{q^{m}-1}{n}}$. Then $\beta$ is a primitive $n$-th root of unity. For $0\leq i\leq n-1$, $m_{i}(x)$ is the minimal polynomial of $\beta^{i}$ over $\mathbb{F}_{q}$. We define the $q$-cyclotomic cosets $C_{i}^{(q,n)}=\{i,iq,\ldots,iq^{l-1}\}{\pmod{n}}$ where $l$ is the smallest positive integer such that $iq^{l}\equiv i{\pmod{n}}$. Then $\{0,1,\ldots,n-1\}$ is partitioned into several $q$-cyclotomic cosets. And
$x^{n}-1={\rm lcm}(m_{0}(x),m_{1}(x),\ldots,m_{n-1}(x))\text{ where }m_{i}(x)=\prod\limits_{j\in C_{i}^{(q,n)}}(x-\beta^{j}).$
Note that we use $i{\pmod{n}}$ to denote the unique integer in the set $\{0,1,\ldots,n-1\}$, which is congruent to $i$ modulo $n$. Thus, we have $m_{i}(x):=m_{i{\pmod{n}}}(x)$.
With these facts, we recall the definition of BCH codes, an important class of cyclic codes.
@@
Let $C_{(q,n,\delta,b)}$ be the cyclic code of length $n$ with the generator polynomial $g_{(q,n,\delta,b)}(x)$. Then $C_{(q,n,\delta,b)}$ is called a BCH code with designed distance $\delta$.
It is well-known that the minimum distance of $C_{(q,n,\delta,b)}$ is greater or equal to $\delta$, see [20].